If women have pulse rates that normally distributed with mean =74 and the standard deviation =12.0, then the probability that one woman randomly selected will have pulse rate less than 68 is approximately?
1
Expert's answer
2020-04-30T17:17:32-0400

"Given \; that, μ=74, σ=12, then,\\
 P(x< 68)=P(Z<\frac{68-74}{12})\\
=P(Z<-0.5)=0.5-P(0<Z<0.5)\\
=0.5-0.1915=0.3085\\"